# Classification and reconstruction of images in the problem single-pixel imaging using classical and quantum neural networks

### v1

@@ Abstract
Single-pixel cameras can be an excellent solution for light ranges outside the visible spectrum, combined with machine learning, they can analyze images quickly enough for practical applications. In the future of the development of quantum technologies, quantum computers can further speed up the solution of such problems. In this work we simulated a single-pixel detection experiment using Hadamard basis patterns, where images from the MNIST handwritten digit dataset were used as objects. There were selected 64 measurements with maximum variance (6% of the number of pixels in the image). We created algorithms for classifying and reconstruction images based on these measurements using classical fully connected neural networks and parameterized quantum circuits. Classical and quantum classifiers showed accuracies of 96% and 95% respectively after 6 training epochs, which is quite competitive result. Image reconstruction was also demonstrated using classical and quantum neural networks after 10 training epochs, the structural similarity index values were 0.76 and 0.25, respectively, which indicates that the problem in such a formulation turned out to be too difficult for quantum neural networks in such a configuration for now.
@@
In single-pixel detection [1], the light intensity scattered by an object is measured by a single photodetector (not matrix as in usual case). Illuminating an object with structured light with a certain set of masks (patterns) allows us to reconstruct an image from such measurements, so that we can obtain spatial resolution without having it in the detector itself. The image is reconstructed by solving the inverse problem, knowing the measurement and pattern at which it was obtained. The single-pixel detection method makes it possible to obtain images at wavelengths outside the visible light range (where is no cheap imaging method, as CCD camera in the visible range), with precise time or depth resolution, and also in turbid environments, which causes greater practical significance of the direction.
Quantum machine learning (QML) is a promising science direction [2, 3, 4, 5], located at the intersection of quantum physics and computer science, in which machine learning methods are developed and studied that can effectively use the unique features of a quantum computer such as superposition and entanglement [6]. In addition, the variational (parameterized) quantum circuits used in quantum machine learning are resistant to the noise of quantum processors [7, 8], which means they can potentially find their useful application before other known quantum algorithms, which require a fault-tolerant quantum computer. Quantum computing is developing to solve various problems in such fields as logistics, finance, medicine and machine learning [9, 10, 11, 12, 13]. This work is aimed at developing this approach as applied to the problem of single-pixel image detection and compare it’s results with classical neural networks (NN), that have shown good results in this area in recent years.
@@ 2 Related work
Single-pixel detection is a promising and cost-effective imaging technique at wavelengths throughout the electromagnetic range (i. g. X-ray [14, 15, 16], terahertz [17, 18, 19]). Single-pixel detection is often used in the near-infrared range due to the availability of detectors with good sensitivity and sources operating in this range. This wavelength range is particularly suitable for detection through scattering media such as fog [20, 21], and has also been used to detect and visualize methane leaks [22].
@@
It is also worth noting that many problems does not require reconstruction of the complete signal, this is the case in applications such as detection or classification [25].
Recently, works on single-pixel imaging using machine learning methods has begun to appear. For example, in [26], using neural networks (NN), a set of patterns that were most effective for an object was constructed, and compression of a set of masks up to 4% was demonstrated to reconstruct 2D images at video signal speed. This model was later used to obtain 3D images in [27]. And in [28], a neural network was used to develop a small number of patterns to classify and identify very fast moving objects. This means that single-pixel cameras with image processing using neural networks are an excellent candidate for many practical applications, such as controlling self-driving cars, night-vision, gas sensing and medical diagnostics. In addition, it is possible to use alternative types of computing such as optical machine learning for image reconstruction [29].
@@ 3 Data
A simple method of obtain an image using a single-pixel detector is to measure each pixel in turn (raster scanning). However, sequentially measuring information about only one pixel in turn is an inefficient use of light capabilities. More common scanning strategy is to use a sequence of spatially resolved light patterns and measure the intensity as different patterns illuminate the object. To reduce the number of measurements required for high quality image restoion a set of orthogonal patterns are used, such as the Hadamard basis.
In this work, a vector of numbers obtained by simulating single-pixel measurements is used as data, a set of Hadamard matrices is used as patterns, and a dataset of images of handwritten digits MNIST (28 28) is used as objects, containing objects of ten classes, images of numbers (from 0 to 9). The dataset consists of a training set of 60 000 images and a test set of 10 000 images.
@@ 4.1 Classical neural networks
For the classic solution, fully connected neural networks were used with as activation function and Adam as optimizer. The metrics used in this work are Cross-Entropy (Eq. 2) loss for classification problem and MSE (Eq. 3) for the regression problem of image reconstruction. The architecture of the classifier consists of one hidden layer ( neurons for layers respectively, that is, 9 610 trainable parameters (according to the Eq. 4)). To reconstruct images, a neural network of four hidden layers is used ( neurons, that is, 20 122 024 trainable parameters (Eq. 4)).
Cross-entropy is given by the following expression:
@@
Mean squared error loss functions can be written as follows:
@@ 4.2 Quantum neural networks
At first we need to encode classical data to qubits, which could be done by the procedure called amplitude embedding so that we prepare our qubits in state , where quantum state complex amplitudes become equals to normalized classical feature values, that allows us to encode features to qubits. It is worth clarifying that this method of data encoding as amplitude embedding is not optimal for near-term hardware installations, since it has a depth that grows exponentially with the number of qubits, but is acceptable for simulation, since it requires a significantly smaller number of qubits than other encoding methods, which is very important for classical simulation. Research into how to most efficiently encode classical data into qubits is now actively underway in the global community [41].
@@ 5 Results
| Network | 111According to Eq. 4 and Eq. 8 for classical and quantum networks respectively | Accuracy | Training time 222Google Colab T4 GPU time for classical solution and PennyLane lightning.qubit simulator on Intel(R) Core(TM) i7-11700 workstation for quantum solution |
| classical | 9 610 | 0.96 | 15.6 s |
@@
The results of training networks to reconstruct images from simulated single-pixel detection experiment are presented in Fig. 5. As you can see the gap between the classical neural network and quantum ones with different numbers of layers is very large. The training time of a quantum neural network depends linearly on the number of layers in it. The dependence of the improvement in the quality of the reconstructed image with an increase in the number of layers does not allow us to hope for the ability to observe a comparable result for a quantum neural network within a reasonable simulation time.
We can also look at how the neural network copes with image reconstruction in a more understandable form by comparing the objects of the experiment we simulated (images from the MNIST dataset) and the images obtained by the neural network, as shown in Fig. 6 and corresponding Structural similarity index metrics (SSIM).
@@ 6 Conclusion
Classical and quantum neural networks were built to solve problems of classification and reconstruction of images based on the measurements of an object based on the MNIST handwritten digits dataset in the Hadamard basis, which is only 6% of the number of pixels in the original object. The constructed classical classifier showed a prediction accuracy of 96%. The developed classical neural network, which reconstructs an image in a single-pixel detection task, made it possible to obtain a high-quality image with a mean-square error of 0.07 and a structural similarity index of 0.76. It was also investigated how the training quality metrics of the created classical neural networks depend on the number of measurements in the Hadamard basis and it was shown that 64 (6% of all possible) for such a task is the optimal value.
The quantum machine learning method was used to solve the problem of classifying “single-pixel images” and thus demonstrated the applicability of quantum neural networks to the problem of analyzing the problem of single-pixel visualization. A quantum neural network was created and trained on a quantum simulator, which predicts the class of a “single-pixel image” with an accuracy of 95%, which is quite competitive result and could be even better with a number of parameters comparable to the classical case. The quantum neural network for image reconstruction has been developed and the results of its work have been demonstrated; it reconstructs images with a structural similarity index of 0.25. Despite the significant difference from images reconstructed using a classical neural network, this result can also be considered successful as proof of concept for now.
Single-pixel detection is a promising and cost-effective imaging technique across the entire electromagnetic spectrum. Combined with recent advances in machine learning algorithms, single-pixel detection promises to be a powerful method for low-cost, scan-free 3D recognition and classification, which holds promise for critical applications such as object detection and classification, surface mapping, and 3D situation recognition for autonomous vehicles. And the results obtained in this work give us hope for new and interesting opportunities that quantum machine learning may provide us in the future.

### v2

# Classification and reconstruction for single-pixel imaging with classical and quantum neural networks

@@ Abstract
Single-pixel cameras are effective solution for imaging outside the visible spectrum where traditional CMOS/CCD cameras have challenges. Combined with machine learning, they can analyze images quickly enough for practical applications. Solving the problem of high-dimensional single-pixel visualization can potentially be accelerated using quantum machine learning, thereby expanding the range of practical problems. In this work we simulated a single-pixel imaging experiment using Hadamard basis patterns, where images from the MNIST handwritten digit dataset were used as objects. There were selected 64 measurements with maximum variance (6% of the number of pixels in the image). We created algorithms for classifying and reconstruction images based on these measurements using classical fully connected neural networks and parameterized quantum circuits. Classical and quantum classifiers showed accuracies of 96% and 95% respectively after 6 training epochs, which is quite competitive result. Image reconstruction was also demonstrated using classical and quantum neural networks after 10 training epochs, the structural similarity index measure values were 0.76 and 0.25, respectively, which indicates that the problem in such a formulation turned out to be too difficult for quantum neural networks in such a configuration for now.
@@
In single-pixel imaging [1], the light intensity scattered by an object is measured by a single photodetector (not matrix as in usual case). Illuminating an object with structured light with a certain set of masks (patterns) allows us to reconstruct an image from such measurements, so that we can obtain spatial resolution without having it in the detector itself. The image is reconstructed by solving the inverse problem, knowing the measurement and pattern at which it was obtained. The single-pixel imaging method makes it possible to obtain images at wavelengths outside the visible light range (where no cheap imaging camera, such as CCD camera, is available), with precise time or depth resolution, and also in turbid environments, which causes greater practical significance of the direction. To analyze “single-pixel images” fast enough for practical applications machine learning is used. Quantum machine learning has the potential to speed up the learning process for large-scale problems, which is of great interest to researchers.
Quantum machine learning (QML) is a promising science direction [2, 3, 4, 5], located at the intersection of quantum physics and computer science, in which machine learning methods are developed and studied that can effectively use the unique features of a quantum computer such as superposition and entanglement [6]. In addition, the variational (parameterized) quantum circuits used in quantum machine learning are resistant to the noise of quantum processors [7, 8], which means they can potentially find their useful application before other known quantum algorithms, which require a fault-tolerant quantum computer. Quantum computing is developing to solve various problems in such fields as logistics, finance, medicine and machine learning [9, 10, 11, 12, 13]. This work is aimed at developing this approach as applied to the problem of single-pixel imaging and compare obtained results with classical neural networks (NN), that have shown good results in this area in recent years.
@@ 2 Related work
Single-pixel imaging is a promising and cost-effective imaging technique at wavelengths throughout the electromagnetic range (i. g. X-ray [14, 15, 16], terahertz [17, 18, 19]). Single-pixel imaging is often used in the near-infrared range due to the availability of detectors with good sensitivity and sources operating in this range. This wavelength range is particularly suitable for imaging through the scattering media such as fog [20, 21], and has also been used to detect and visualize methane leaks [22].
@@
It is also worth noting that many problems does not require reconstruction of the original object, this is the case in applications such as detection or classification [25].
Recently, works on single-pixel imaging using machine learning methods has begun to appear. For example, in [26], using neural networks (NN), a set of patterns that were most effective for an object was constructed, and compression of a set of masks up to 4% was demonstrated to reconstruct 2D images at video signal speed. This model was later used to obtain 3D images in [27]. In [28], a neural network was used to develop a small number of patterns to classify and identify very fast moving objects. Thus, the single-pixel cameras with image processing using neural networks are an excellent candidate for many practical applications, such as controlling self-driving cars, night-vision, gas sensing and medical diagnostics. In addition, it is possible to use alternative types of computing such as optical machine learning for image reconstruction [29].
@@ 3 Data
A simple method to obtain an image using a single-pixel detector is to measure each pixel in turn (raster scanning). However, sequentially measuring information about only one pixel in turn is an inefficient use of light capabilities. More common scanning strategy is to use a sequence of spatially resolved light patterns and measure the intensity as different patterns illuminate the object. To reduce the number of measurements required for high quality image reconstruction a set of orthogonal patterns are used, such as the Hadamard basis.
In this work, a vector of numbers obtained by simulating single-pixel measurements is used as data, a set of Hadamard matrices is used as patterns, and a dataset of images of handwritten digits MNIST (28 28, numbers from 0 to 9) is used as objects divided to ten classes. The dataset consists of a training set of 60 000 images and a test set of 10 000 images.
@@ 4.1 Classical neural networks
For the classic solution fully connected neural networks were used, with as the activation function, optimized by Adam. The metrics used in this work are Cross-Entropy (Eq. 2) loss for image classification problem and Mean squared error (MSE, Eq. 3) for the regression problem of image reconstruction. The architecture of the classifier consists of one hidden layer ( neurons for layers respectively, that is, 9 610 trainable parameters (according to the Eq. 4)). To reconstruct images, a neural network of four hidden layers is used ( neurons, that is, 20 122 024 trainable parameters (Eq. 4)).
The cross-entropy loss function is given by the following expression:
@@
The MSE loss function can be written as follows:
@@ 4.2 Quantum neural networks
At first we need to encode classical data to qubits, which could be done by the procedure called amplitude embedding so that we prepare our qubits in state , where quantum state complex amplitudes become equals to normalized classical feature values, which are real values. That allows us to encode features to qubits. It is worth clarifying that this method of data encoding as amplitude embedding is not optimal for near-term hardware installations, since the circuit for amplitude embedding has a depth that grows exponentially with the number of qubits, but is acceptable for simulation, since it requires a significantly smaller number of qubits than other encoding methods, which is very important for classical simulation. Research into how to most efficiently encode classical data into qubits is now actively underway in the global community [41].
@@ 5 Results
| Network | 111According to Eq. 4 and Eq. 8 for classical and quantum networks respectively | Accuracy | Training time 222CPU time for classical solution and PennyLane lightning.qubit simulator on CPU for quantum solution (CPU: Intel(R) Core(TM) i7-11700 workstation) |
| classical | 9 610 | 0.96 | 9.3 s |
@@
The results of training networks to reconstruct images from simulated single-pixel imaging experiment are presented in Fig. 5. As you can see the gap between the classical neural network and quantum ones with different numbers of layers is very large. The training time of a quantum neural network depends linearly on the number of layers in it. The dependence of the improvement in the quality of the reconstructed image with an increase in the number of layers does not allow us to hope for the ability to observe a comparable result for a quantum neural network within a reasonable simulation time.
We can also look at how the neural network copes with image reconstruction in a more understandable form by comparing the objects of the experiment we simulated (images from the MNIST dataset) and the images obtained by the neural network, as shown in Fig. 6 and corresponding Structural similarity index measure (SSIM).
@@ 6 Conclusion
Classical and quantum neural networks were built to solve problems of classification and reconstruction of images based on the measurements of an object founded on the MNIST handwritten digits dataset in the Hadamard basis, which is only 6% of the number of pixels in the original object. The constructed classical classifier showed a prediction accuracy of 96%. The developed classical neural network, which reconstructs an image in a single-pixel imaging task, made it possible to obtain a high-quality image with a mean-square error of 0.07 and a structural similarity index measure of 0.76. It was also investigated how the training quality metrics of the created classical neural networks depend on the number of measurements in the Hadamard basis and it was shown that 64 (6% of all possible) for such a task is the optimal value.
The quantum machine learning method was used to solve the problem of classifying “single-pixel images” and thus demonstrated the applicability of quantum neural networks to the problem of analyzing the problem of single-pixel visualization. A quantum neural network was created and trained on a quantum simulator, which predicts the class of a “single-pixel image” with an accuracy of 95%, which is quite competitive result and could be even better with a number of parameters comparable to the classical case. The quantum neural network for image reconstruction has been developed and the results of its work have been demonstrated; it reconstructs images with a structural similarity index measure of 0.25. Despite the significant difference from images reconstructed using a classical neural network, this result can also be considered successful as proof of concept for now.
Single-pixel imaging is a promising and cost-effective imaging technique across the entire electromagnetic spectrum. Combined with recent advances in machine learning algorithms, single-pixel imaging promises to be a powerful method for low-cost, scan-free 3D recognition and classification, which holds promise for critical applications such as object detection and classification, surface mapping, and 3D situation recognition for autonomous vehicles. And the results obtained in this work give us hope for new and interesting opportunities that quantum machine learning may provide us in the future.